The sum of Jane's age and Jim's age is 40 years. Jane's age 10 years from now will be 1 year less than 4 times Jim's age 6 years ago. Find their ages.
Answer by gsmani_iyer(201)


Let the age of Jane = x;
Let the age of Jim = y;
From the problem, we will frame equations.
x + y = 40 ....... (1) The sum of the ages.

x + 10 = 4(y-6) - 1 .... (2) The age of Jane after 10 years.

So, x + 10 = 4y-24-1 .... (2) is expanded.

x + 10 = 4y - 25 ..... (3)

i.e x - 4y = -10 -25 = x - 4y = -35 ....(4)

(1) - (4), we get

5y = 75; So y = 75/5 = 5.

By substituting this value in Eq.(1)

x + 5 = 40; i.e. x = 40 - 5 = 35

x = 35; y = 5. Answer.

I would advise you to check my earlier solved linear problems so that you can solve such problems yourself.

gsm.